OpenStudy (anonymous):

Find the slope, if possible, of the line passing through each pair of points. (2, -1) and (6,3)

OpenStudy (anonymous):

slope = y2-y1/ (x2-x1) slope = 3- (-1) /(6-2) = 4/4 = 1